OpenStudy (anonymous):

y = 3x + 1 a. (1, -4) b. (-2, -5)

OpenStudy (akashdeepdeb):

Just put the value of the points AND SEE which one fits the equation! That means LHS should be equal to RHS!! :D

OpenStudy (anonymous):

@AkashdeepDeb I tried doing that but I got wrong answers so thats why im here so someone can show me to do it right

OpenStudy (akashdeepdeb):

Try it with the B) option y = 3x + 1 point is (-2,-5) RHS = 3(-2) + 1 = -6+1 = -5 LHS = y = -5 LHS=RHS So (-2,-5) IS a point on the line!! :D Understood? :)

OpenStudy (anonymous):

@AkashdeepDeb Yea it makes sense now, thank!

OpenStudy (akashdeepdeb):

:)